can someone please help me with this problem:
How many 3-digit numbers can be formed using the digits 0,1,2,3,4,5,6 if repetition of digits is not allowed?
a. 210
b. 5
c. 6
d. 343

Answer by jim_thompson5910(28598)

You have 7 digits to choose from. After using one of them, you will only have 6 and after using another, you will only have 5. So this means that the number of possible combinations are


7*6*5 = 42*5 = 210

So there are 210 ways that a three digit number can be formed using the digits 0,1,2,3,4,5,6 and repetition is not allowed